OpenStudy (anonymous):

y = -2 x + -2 how would you graph this?

OpenStudy (happykiddo):



OpenStudy (mathstudent55):

This line is in the slope-intercept form which is in general y = mx + b, where m = slope and b = y-intercept. This means the following: The y-intercept tells you where the line crosses the y-axis. Since b = -2, that means it crosses the y-axis at y = 2, which is the point (0, 2) Start a graph, and plot the point (0, 2) The the m, which is the slpoe, tells you how the line tilts. Since m = -2 and the slope = rise/run, or the amount of change in y divided by the amount of change in x, think of m = -2 as m = -2/1 = 2/-1 That means after you start with the point (0, -2), move down 2 (-2) and 1 to the right. That's another point. Now that you have two points, you can draw a line through them and extend it.

OpenStudy (happykiddo):

my drawing is kinda crappy but mathstudent55 gives a good explanation

OpenStudy (anonymous):

They were both very helpful. Thank you! :)